# Debiasing astro-Photometric Observations with Corrections Using Statistics (DePhOCUS)

### v1

@@ Abstract
Photometric measurements allow the determination of an asteroid’s absolute magnitude, which often represents the sole means to infer its size. Photometric observations can be obtained in a variety of filters that can be unique to a specific observatory. Those observations are then calibrated into specific bands with respect to reference star catalogs. In order to combine all the different measurements for evaluation, photometric observations need to be converted to a common band, typically V-band. Current band-correction schemes in use by IAU’s Minor Planet Center (MPC), JPL’s Center for Near Earth Object Studies (CNEOS) and ESA’s NEO Coordination Centre (NEOCC) use average correction values for the apparent magnitude derived from photometry of asteroids as the corrections are dependent on the typically unknown spectrum of the object to be corrected. By statistically analyzing the photometric residuals of asteroids, we develop a new photometric correction scheme that does not only consider the band, but also accounts for reference catalog and observatory. We describe a new statistical photometry correction scheme for asteroid observations with debiased corrections. Testing this scheme on a reference group of asteroids, we see a $36\%$ reduction in the photometric residuals. Moreover, the new scheme leads to a more accurate and debiased determination of the $H$-$G$ magnitude system and, in turn, to more reliable inferred sizes. We discuss the significant shift in the corrections with this “DePhOCUS” debiasing system, its limitations, and the impact for photometric and physical properties of all asteroids, especially Near-Earth Objects.
@@ 1 Introduction
The physical description of the asteroid’s brightness depends on two parameters in the $H$-$G$ model by Bowell et al. (1989): $H$ is the absolute magnitude and $G$ is the slope parameter, describing the steepness of the $H$-$G$ phase function with increasing solar phase angle (cf. Fig 1 and Eq. 3 in the section 2.1-2.2). Even though this method is well established, the procedure has considerable uncertainties and limitations:
@@
All of this shows that determining the phase curve of an asteroid with measuring the asteroid’s rotational lightcurve, covering a wide range of phase angles, and achieving highly accurate (and calibrated) photometry (Buchheim, 2010) poses major difficulties for astronomers and surveys that primarily are focused on astrometry.
Moreover, the $H$-$G$ model itself has some intrinsic limitations. The so-called opposition effect and the light back-scattering in general, especially for high- and low-albedo asteroids, is not well considered by the model (Muinonen et al., 2002, 2010). The International Astronomical Union (IAU) officially adopted the updated three-parameter $H,G_{1},G_{2}$ model by Muinonen et al. (2010) in 2012. With an updated fitting algorithm and a two-parameter version ($H,G_{12}^{*}$) by Penttilä et al. (2016), the $H$-$G$ model by Bowell et al. (1989) is theoretically replaced by now. But still, as most phase curves are too inaccurate to determine even one of the $G$ values, the $H$-$G$ model is still the most frequently used model when dealing with astro-photometry only datasets.
@@
Large surveys like the Panoramic Survey Telescope And Rapid Response System (Pan-STARRS), Asteroid Terrestrial-impact Last Alert System (ATLAS) and the Zwicky Transient Facility (ZTF) were used to derive phase curves from asteroids with their own astro-photometry, trying to infer the slope parameter in order to classify taxonomic types and to correlate with the albedo (Vereš et al., 2015; Mahlke et al., 2021; Carry et al., 2024). However, when using different photometric systems and converting to V-band, potentially inaccurate color corrections are used, potentially introducing systematic errors.
Colazo et al. (2021) used measurements of asteroids also from Gaia in the Gaia DR2 data-release to determine the $H$-$G$ phase function parameters with the instrument’s $g$ magnitude for nearly 14 000 objects. As the study also included ground-based observations in order to have near-opposition phase angle data, the issue of combining different photometric color-band data again becomes evident.
@@
The MPC and others tried to establish observatory-dependent corrections to the reported apparent magnitudes from observers (Vereš et al., 2015), but so far none of these has implemented such methods. We are going to quantify observatory-dependent biases in astro-photometry and implement a correction algorithm of these biases in order to reduce statistical uncertainties and systematical errors in the $H$, $G$ parameter fitting.
@@ 2 Methods
Observational data submitted to the MPC by observers throughout the world is used for the analysis. We use all observations with valid photometric measurements, i.e. apparent magnitudes in a specified color band. These values are first pre-processed, stored and then analyzed in comparison to an accurate photometric reference source. With that we are able to compute the photometric residuals of all observations from different observatory stations. The residuals undergo a statistical analysis to determine possible significant deviations and uncertainties. The results for the systematic biases will then be implemented in a correction algorithm.
@@ 2.1 Data sources and pre-processing pipeline
in order to normalize the effects of distances on the brightness for asteroids (Williams, 2013). The reduced magnitude is then used to visualize the asteroid’s phase curve ($m_{red}$ over $\alpha$, cf. Fig. 1), separated for the different color bands that are used in the observations. Additionally, plots of the range of phase angles covered in the observations can be generated for the asteroid, alongside with a data file that summarizes all the obtained values in the pre-processing pipeline.
Figure 1: $H$-$G$ phase function for an asteroid with $H=10$ mag and varying $G$.
@@ 2.2 Data analysis system
For a given dataset of validated asteroids the observations from all observatory sites are statistically analyzed for biases. Therefore, we have a look at the photometric residuals of the observations
@@
The fundamental parameters for this magnitude system, the absolute magnitude $H$ and slope parameter $G$, are extracted from the prepublished database from the Ondrejov Asteroid Photometry Project updated as of 2023-08-11333https://www.asu.cas.cz/~ppravec/newres.htm by Pravec et al. (2012). This database contains results from asteroid rotational and binary light curve measurements, among others the $H$-$G$ parameters, for more than 1000 asteroids. The methodology of the measurements and their analysis is described by Pravec et al. (1996, 1997). Their derived asteroid absolute magnitudes $H$ correspond to the mean light (effectively an average over asteroid’s rotation) in the Johnson V photometric system, calibrated with the Landolt (1992) standard stars.
The data analysis system as displayed in Fig. 3 uses as a standard reference all asteroids in the database with $H$-$G$ magnitude system parameters in the V-band. In case of multiple measurements for the same asteroid the newer updated parameters are used. With that, we get a total of 468 different asteroids with reference $H$-$G$ model parameters. For all these asteroids, all available observations at MPC are downloaded and pre-processed. Using the $H$ and $G$ parameter from Pravec et al. (2012) in the $H$-$G$ phase function with the corresponding phase angle of the observation and comparing it to the more than 450 000 reduced magnitude measurements, we get a database with all the residuals of observations from different observatories (up to 2023-08-11). This database not only includes the residuals, but also raw data (e.g., apparent magnitude measurement) and meta-data (e.g., color band, star catalog, observatory station, time).
@@
There might be an additional influence of the spectra of the individual asteroids, which would lead to deviating results. A high number of observations for these objects would dominate in the data sample and introduce a bias. Therefore, the statistics should be independent of the varying number of observations used from each asteroid. For this reason, multiple observations of the same object in the same group (B, BC or BCO) should be summarized into one dataset by averaging the values. This can be achieved with the existing method by using the asteroid name as an additional grouping parameter and performing the analysis as for the previous cases. These additional asteroid-dependent analyses are correspondingly named aB, aBC and aBCO analysis. Before exporting specific biases and their errors, an average across all the different asteroids within the B, BC or BCO group is calculated. With this kind of analysis effectively all asteroids used in one group are getting weighted equally, thus minimizing the effect of intrinsic photometric properties of individual asteroids, especially in cases with a strongly deviating amount of observations per asteroid. The data analysis system hereby provides a statement about the bias for the average asteroid.
@@ 2.3.1 Current correction systems
Table 1: Current astro-photometric correction offsets (in mag) to a reference V-band used by NEOCC (until 2023-09-28) [^5] Photometry correction used by NEOCC have been updated on 2023-09-28 based on the results of this study (https://neo.ssa.esa.int/system-status)., MPC and JPL in their orbital determination systems (Vereš et al., 2018). Values are sorted by the color band’s effective wavelength $\lambda_{\text{eff}}$ and its bandwidth $\Delta\lambda_{\text{eff}}$.
@@ 2.3.2 Statistical correction algorithm
We are using the results from the bias analysis system in Sec. 2.2 to develop an iterative algorithm that corrects the specific systematic biases in the apparent magnitude measurements for NEOs. For this, the average bias values $\overline{\delta_{m}}$ obtained by the aB, aBC and aBCO analyses are used as an offset for the correction, which needs to be added on the observational apparent magnitude (cf. definition of the residuals $\delta m_{red}$ in Eq. (2)).
@@
with the number $n$ of observations residuals $\delta_{m}$ in the new correction. A two-tailed test is then performed with the cumulative distribution function of the Student’s t-distribution $F(t;\nu)$, with $\nu=n-1$ the degree of freedom, by checking if the two-tailed probability $2F(t;\nu)$ is larger than a given $p$-level. If this is the case, the new correction is deviating significantly and therefore gets appended to the list of correction rules. By repeating this algorithm for all possible corrections step-wise, the final correction system is defined, with all corrections having a significance larger than a given $p$-level.
@@ 2.3.3 Light curve effects
Since most astrometric observers only make a few observations of an object during the night, with rough estimated astro-photometry, the observational coverage does not account for light curve effects that are due to rotation and the shape of the asteroid, which typically happens on a period longer than the duration of the observation for main-belt asteroids – for NEOs rotation periods can also be as short as $2.5$ seconds (Devogèle et al., 2024; Thirouin et al., 2016). Overall, these effects create a periodic variation in brightness, thus introducing natural deviations from the $H$-$G$ system phase curve.
@@
Assuming the variation of light curve has a sinusoidal shape, the expected RMS of the observations due to the rotational effects is given by $1/2\sqrt{2}$ times the peak-to-trough amplitude, which leads us to an average effect on the RMS for asteroids of $(0.1417-0.1429)$ mag. Including the typical numerical precision of astro-photometrical observations of $0.10$ mag, a minimum RMS of roughly $0.25$ mag is expected as a physical boundary, if we do not consider and compensate for light curve effects. As shown in other studies, this value is also assumed to be a typical error for the $H$ parameter (Mainzer et al., 2012).
@@ 3.1 Bias analysis
We use the methods from Sec. 2.2 to analyze the distribution of specific biases in color bands (3.1.1), catalogs (3.1.2) and observatories (3.1.3). From the Ondrejov Asteroid Photometry Project database 394 out of the 468 asteroids ($\approx 84\%$) with V-band measurements for the H and G parameter were used for the analysis. These are selected by having an export date of the measurements earlier than 2021-10-21. The more recent measurements will be used in Sec. 3.3 for the validation of the correction system.
@@ 3.1.1 Band bias results
In Fig. 6 the obtained biases are also represented in a graphical way by showing the distribution of the bias values in each color band. Most of them have symmetrical peak-shapes, with some having narrow widths like g, c, w, o or i and others having broader widths like V, R, G or ‘blank’, where the latter are the ones with noticeable deviations to the current correction, as described before. Most interesting are the color bands where there are asymmetric or double peaked shapes, like the r- and I-band. In these cases, one portion of the observations is having one specific bias, conspicuously around 00 mag, and one portion is having another bias, around the value of the correction which would be expected. A bias of 00 mag means, that there is no bias and no correction is needed. As there should be bias expected by the intrinsic color properties of the light reflected by the asteroid as derived by Williams (2013), it can be suspected that there are observers or measurers who provide the wrong color band in their reports, or even already correct their photometry to V with the correction values by themselves (while still reporting the original band in the MPC submission). Those would be corrected a second time by the MPC, thus aggravating the results. This underlines the importance of investigating the biases of individual observatories.
@@ 3.1.2 Catalog bias results
Binning the observations in each color band by the reference catalog used, we can analyze systematic magnitude biases by catalogs, independent of the color band bias. The catalog refers to the astrometric reference star catalog used to derive the measurements, which is encoded by a single character [^9] https://www.minorplanetcenter.net/iau/info/CatalogueCodes.html in column 72 of the observations in the MPC 80-column format, which has been introduced by Williams (2013). More recent observations with the Astrometry Data Exchange Standard (ADES) would allow the report of a separate photometric catalog, but as this option was not that extensively used until very recently, we assume that there is no major deviation between astrometric and photometric reference star catalog.
@@
(continued on next page)
Figure 8: Results from aBC bias analysis with data from 394 out of 468 asteroids, with the average bias and its standard error in the different color bands and astrometric catalogs, sorted in decreasing order of the number of observations, shown per band with comparison band bias (colored line) including error margins (blue area) from the aB analysis (Tab. 2), leaving out the following bands with just one catalog: y, Y, H, K (only 2MASS) and L (only Gaia DR2)
We can see in the analysis that there is a significant influence of the reference star catalog used on the bias and accuracy of the photometry. In some color bands like V, R, G and ‘blank’ the results are varying by up $0.5$ mag, with only a few catalogs lying in the uncertainty margins of the overall color band bias. For the ‘blank’-band this behavior is expected as the observations might be taken in fact in various color bands that correspond with certain catalogs (e.g., Gaia catalogs with the Gaia G-band), which were just not provided in the report.

### v2

@@ Abstract
Photometric measurements allow the determination of an asteroid’s absolute magnitude, which often represents the sole means to infer its size. Photometric observations can be obtained in a variety of filters that can be unique to a specific observatory. Those observations are then calibrated into specific bands with respect to reference star catalogs. In order to combine all the different measurements for evaluation, photometric observations need to be converted to a common band, typically V-band. Current band-correction schemes in use by IAU’s Minor Planet Center (MPC), JPL’s Center for Near Earth Object Studies (CNEOS) and ESA’s NEO Coordination Centre (NEOCC) use average correction values for the apparent magnitude derived from photometry of asteroids as the corrections are dependent on the typically unknown spectrum of the object to be corrected. By statistically analyzing the photometric residuals of asteroids, we develop a new photometric correction scheme that does not only consider the band, but also accounts for reference catalog and observatory. We analyzed nearly 500 000 observations submitted to the MPC from 468 asteroids with published and independently determined high confidence $H$ and $G$ values. We describe a new statistical photometry correction scheme for asteroid observations with debiased corrections. Testing this scheme on a reference group of asteroids, we see a $36\%$ reduction in the photometric residuals. Moreover, the new scheme leads to a more accurate and debiased determination of the $H$-$G$ magnitude system and, in turn, to more reliable inferred sizes. We discuss the significant shift in the corrections with this “DePhOCUS” debiasing system, its limitations, and the impact for photometric and physical properties of all asteroids, especially Near-Earth Objects.
@@ 1 Introduction
The physical description of the asteroid’s brightness depends on two parameters in the $H$-$G$ model by Bowell et al. (1989): $H$ is the absolute magnitude and $G$ is the slope parameter, describing the steepness of the $H$-$G$ phase function with increasing solar phase angle (cf. Fig 1 and Eq. 3 in the section 2.1-2.2). Even though this method is well established, the model, its fitting procedure and results have considerable uncertainties and limitations:
@@
All of this shows that determining the phase curve of an asteroid, achieving highly accurate (and calibrated) photometry (Buchheim, 2010), whilst including the asteroid’s rotational light curve and covering a wide range of phase angles, poses major difficulties for astronomers and surveys that primarily are focused on astrometry.
Moreover, the $H$-$G$ model itself has some intrinsic limitations. The so-called opposition effect and the light back-scattering in general, especially for high- and low-albedo asteroids (Harris and Young, 1989; Harris et al., 1989; Shevchenko et al., 1997; Belskaya and Shevchenko, 2000), is not well considered by the model (Muinonen et al., 2002, 2010). The International Astronomical Union (IAU) officially adopted the updated three-parameter $H,G_{1},G_{2}$ model by Muinonen et al. (2010) in 2012. With an updated fitting algorithm and a two-parameter version ($H,G_{12}^{*}$) by Penttilä et al. (2016), the $H$-$G$ model by Bowell et al. (1989) is theoretically replaced by now. But still, as most phase curves are too inaccurate to determine even one of the $G$ values, the $H$-$G$ model is still the most frequently used model when dealing with astro-photometry only datasets.
@@
Large surveys like the Panoramic Survey Telescope And Rapid Response System (Pan-STARRS) (Kaiser et al., 2010), Asteroid Terrestrial-impact Last Alert System (ATLAS) (Tonry et al., 2018) and the Zwicky Transient Facility (ZTF) (Bellm et al., 2019) were used to derive phase curves from asteroids with their own astro-photometry, trying to infer the slope parameter in order to classify taxonomic types and to correlate with the albedo (Vereš et al., 2015; Mahlke et al., 2021; Carry et al., 2024).
Colazo et al. (2021) used measurements of asteroids also from Gaia in the Gaia DR2 data-release to determine the $H$-$G$ phase function parameters with the instrument’s $g$ magnitude for nearly 14 000 objects. As the study also included ground-based observations in order to have near-opposition phase angle data, different photometric systems needed to be used and converted. When combining multiple color-band data, a conversion to the V-band with color corrections is typically used. Inaccurate color corrections might introduce systematic errors.
@@
The MPC and others demonstrated the need to establish observatory-dependent corrections to the reported apparent magnitudes from observers (Jurić et al., 2002; Vereš et al., 2015), but these methods have not yet been implemented to correct photometry from all of the more than 2500 observatories registered with the MPC. We are now going to quantify the required observatory-dependent biases in astro-photometry and implement a correction algorithm of these biases in order to reduce statistical uncertainties and systematical errors in the $H$, $G$ parameter fitting.
@@ 2 Methods
Observational data submitted to the MPC by observers throughout the world is used for the analysis. We use all observations with valid photometric measurements, i.e. apparent magnitudes in a specified color band. These values are first pre-processed, stored and then analyzed in comparison to an accurate photometric reference source containing 468 asteroids. With that we are able to compute the photometric residuals of all observations from different observatory stations. The residuals undergo a statistical analysis to determine possible significant deviations and uncertainties. The results for the systematic biases will then be implemented in a correction algorithm.
@@ 2.1 Data sources and pre-processing pipeline
in order to normalize the effects of distances on the brightness for asteroids (Williams, 2013). The reduced magnitude is then used to visualize the asteroid’s phase curve ($m_{red}$ over $\alpha$), separated for the different color bands that are used in the observations (cf. Fig. 1). Additionally, plots of the range of phase angles covered in the observations can be generated for the asteroid, alongside with a data file that summarizes all the obtained values in the pre-processing pipeline.
Figure 1: $H$-$G$ phase function for asteroid (99942) Apophis ($H=18.945$, $G=0.15$; from NEOCC) with all available observations at MPC in different color-bands, including the residuals below.
@@ 2.2 Data analysis system
For a given dataset of accurately photometrically validated asteroids the observations from all observatory sites are statistically analyzed for biases. Therefore, we inspect the photometric residuals of the observations
@@
The fundamental parameters for this magnitude system, the absolute magnitude $H$ and slope parameter $G$, are extracted from the prepublished database from the Ondrejov Asteroid Photometry Project updated as of 2023-08-11333https://www.asu.cas.cz/~ppravec/newres.htm by Pravec et al. (2012). This database contains results from asteroid rotational and binary light curve measurements, among others the $H$-$G$ parameters, for more than 1000 asteroids. The methodology of the measurements and their analysis is described by Pravec et al. (1996, 1997). Their derived asteroid absolute magnitudes $H$ correspond to the mean light (effectively an average over asteroid’s rotation) in the Johnson V photometric system, calibrated with the Landolt (1992) standard stars. Using calibrated measurements that consider the asteroids’ light curves minimizes photometric errors in the $H$-$G$ parameters. Despite having a comparably low quantity of measurements, we consider this method to be effective in order to avoid intrinsic biases.
The data analysis system as displayed in Fig. 3 uses all asteroids in the database with $H$-$G$ magnitude system parameters in the V-band as a standard reference. In case of multiple measurements for the same asteroid the newer updated parameters are used. With that, we get a total of 468 different asteroids, including 327 NEOs, with their reference $H$-$G$ model parameters. For all these asteroids, all available observations from the MPC are downloaded and pre-processed. The number of observations per object ranges from 51 up to 7059, with a median of 532 (cf. distribution in Fig. 13 in the appendix). Each object has observations in at least 3 different color bands (maximum 19), with an median of 10 bands available. Using the $H$ and $G$ parameter from Pravec et al. (2012) and the corresponding phase angles of the observations we can compute the $H$-$G$ phase function values. Comparing those values to the more than 450 000 reduced magnitude measurements from observations, we get a database with observational residuals from all different observatories (up to 2023-08-11). This database not only includes the residuals, but also raw data (e.g., apparent magnitude measurement) and meta-data (e.g., color band, star catalog, observatory station, time).
@@
There might be an additional influence of the spectra of the individual asteroids, which would lead to deviating results. A high number of observations for these objects would dominate in the data sample and introduce a bias. Therefore, the statistics should be independent of the varying number of observations used from each asteroid (cf. Fig 13). For this reason, multiple observations of the same object in the same group (B, BC or BCO) should be summarized into one dataset by averaging the values. This can be achieved with the existing method by using the asteroid name as an additional grouping parameter and performing the analysis as for the previous cases. These additional asteroid-dependent analyses are correspondingly named aB, aBC and aBCO analysis. Before exporting specific biases and their errors, an average across all the different asteroids within the B, BC or BCO group is calculated. With this kind of analysis effectively all asteroids used in one group are getting weighted equally, thus minimizing the effect of intrinsic photometric properties of individual asteroids, especially in cases with a strongly deviating amount of observations per asteroid. The data analysis system hereby provides a statement about the bias for the average asteroid.
@@ 2.3.1 Current correction systems
Table 1: Current astro-photometric correction offsets (in mag) to a reference V-band used by NEOCC (before 2023-09-28) [^5] Photometry correction used by NEOCC have been updated on 2023-09-28 based on the results of this study (https://neo.ssa.esa.int/system-status)., MPC and JPL in their orbital determination systems (Vereš et al., 2018). Values are sorted by the color band’s effective wavelength $\lambda_{\text{eff}}$ and its bandwidth $\Delta\lambda_{\text{eff}}$.
@@ 2.3.2 Statistical correction algorithm
We are using the results from the bias analysis system in Sec. 2.2 to develop an iterative algorithm that corrects the specific systematic biases in the apparent magnitude measurements for asteroids and especially NEOs, as well. For this, the average bias values $\overline{\delta_{m}}$ obtained by the aB, aBC and aBCO analyses are used as an offset for the correction, which needs to be added on the observational apparent magnitude (cf. definition of the residuals $\delta m_{red}$ in Eq. (2)).
@@
with the number $n$ of observations residuals $\delta_{m}$ in the new correction. A two-tailed test is then performed with the cumulative distribution function of the Student’s t-distribution $F(t;\nu)$, with $\nu=n-1$ the degree of freedom, by checking if the two-tailed probability $p=2F(t;\nu)$ is larger than or equal to a given significance level $\alpha$ (null hypothesis $H_{0}$: $p\geq\alpha$. If this is not the case, the new correction is deviating significantly (rejection of the null hypothesis $H_{1}$: $p<\alpha$) and therefore gets appended to the list of correction rules. By repeating this algorithm for all possible corrections step-wise, the final correction system is defined, with all corrections having a confidence $\gamma=1-\alpha$ larger than a given level.
@@ 2.3.3 Light curve effects
Since most astrometric observers only make a few observations of an object during the night, with rough estimated astro-photometry, the observational coverage does not account for light curve effects that are due to rotation and the shape of the asteroid, which typically happens on a period longer than the duration of the observation for main-belt asteroids – for NEOs rotation periods can significantly shorter, with some already observed as brief as $2.5$ seconds (Devogèle et al., 2024; Thirouin et al., 2016). Overall, these effects create a periodic variation in brightness, thus introducing natural deviations from the $H$-$G$ system phase curve.
@@
Assuming the variation of light curve has a sinusoidal shape, the expected RMS of the observations due to the rotational effects is given by $1/2\sqrt{2}$ times the peak-to-trough amplitude, which leads us to an average effect on the RMS for asteroids of $(0.1417-0.1429)$ mag. Including the typical numerical precision of astro-photometrical observations of $0.10$ mag, a minimum RMS of roughly $0.25$ mag is expected as a physical boundary, if we do not consider and compensate for light curve effects. [^9] Also other factors influence the errors in the $H$ and $G$ parameters, e.g. the phase angle coverage, leading to a further increased expected RMS. As shown in other studies, this value is also assumed to be a typical error for the $H$ parameter (Mainzer et al., 2012).
@@ 3.1 Bias analysis
We use the methods from Sec. 2.2 to analyze the distribution of specific biases in color bands (3.1.1), catalogs (3.1.2) and observatories (3.1.3). From the Ondrejov Asteroid Photometry Project database 394 out of the 468 asteroids ($\approx 84\%$) with V-band measurements for the H and G parameter were used for the analysis. These are selected by having an export date of the measurements earlier than 2021-10-21. The subset of the remaining 74 asteroids with more recent measurements will be used for the validation of the correction system (Sec. 3.3).
@@ 3.1.1 Band bias results
In Fig. 6 the obtained biases are also represented in a graphical way by showing the distribution of the bias values in each color band. Most of them have symmetrical peak-shapes, with some having narrow widths like g, c, w, o or i and others having broader widths like V, R, G or ‘blank’, where the latter are the ones with noticeable deviations to the current correction, as described before. Most interesting are the color bands where there are asymmetric or double peaked shapes, like the r- and I-band. A bi-modality test by Hartigan and Hartigan (1985); Hartigan (1985) and implemented by diptest [^10] https://github.com/RUrlus/diptest/releases/tag/v0.8.1 shows a $p$-value of the dip statistics of 0.0086 for r-band and 0.9822 for I-band, meaning r-band is most likely having a bi-modality. In both cases, one portion of the observations is having one specific bias, conspicuously around 00 mag, and one portion is having another bias, around the value of the correction which would be expected. A bias of 00 mag means, that there is no bias and no correction is needed. As there should be bias expected by the intrinsic color properties of the light reflected by the asteroid as derived by Williams (2013), it can be suspected that there are observers or measurers who provide the wrong color band in their reports, or even already correct their photometry to V with the correction values by themselves (while still reporting the original band in the MPC submission). Those would be corrected a second time by the MPC, thus aggravating the results. This underlines the importance of investigating the biases of individual observatories.
@@ 3.1.2 Catalog bias results
(continued on next page)
Figure 8: Results from aBC bias analysis with data from 394 out of 468 asteroids, with the average bias and its standard error in the different color bands and astrometric catalogs (one letter format [^11] https://www.minorplanetcenter.net/iau/info/CatalogueCodes.htmlin brackets), sorted in decreasing order of the number of observations, shown per band with comparison band bias (colored line) including error margins (blue area) from the aB analysis (Tab. 2), leaving out the following bands with just one catalog: y, Y, H, K (only 2MASS) and L (only Gaia DR2)
Binning the observations in each color band by the reference catalog used, we can analyze systematic magnitude biases by catalogs, independent of the color band bias. The catalog refers to the astrometric reference star catalog used to derive the measurements, which is encoded by a single character in column 72 of the observations in the MPC 80-column format, which has been introduced by Williams (2013). More recent observations with the Astrometry Data Exchange Standard (ADES) would allow the report of a separate photometric catalog, but as this option was not that extensively used until very recently, we assume that there is no major deviation between astrometric and photometric reference star catalog.
@@
We can see in the analysis that there is a significant influence of the reference star catalog used on the bias and accuracy of the photometry. In some color bands like V, R, G and when no band is specified (‘blank’) the results are varying by up $0.5$ mag, with only a few catalogs lying in the uncertainty margins of the overall color band bias. For the ‘blank’-band this behavior is expected as the observations might be taken in fact in various color bands that correspond with certain catalogs (e.g., Gaia catalogs with the Gaia G-band), which were just not provided in the report.